# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ 6.1 Proof of Lemma 9
W.l.o.g, $H$ has at least two components; if $H$ consists of a single component, it has to be huge as $\lvert V(H)\rvert=\lvert V(G)\rvert\geq\frac{4}{\varepsilon}\geq 10$. Let $A$ be a node in ^$\bm{\hat{}}{G}_{H}$. Let $K$ be some cycle in ^$\bm{\hat{}}{G}_{H}$ containing $A$ which must exist (and easily found) as ^$\bm{\hat{}}{G}_{H}$ is 2EC and has at least 2 nodes.
@@
Otherwise, arguing as before, we can find another cycle $K^{\prime}$ through $B$ in ^$\bm{\hat{}}{G}_{H_{1}}$ and add $K^{\prime}$ to $H_{1}$ to obtain $H_{2}$, i.e., $H_{2}\coloneq H\cup K\cup K^{\prime}$. Observe that $H_{2}$ is again a canonical bridgeless 2-edge cover and the newly formed component (containing the vertices of $B$) has at least 12 vertices, and is huge. We have $|H_{2}|=|H|+|K|+|K^{\prime}|$ and $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H_{1})-(|K^{\prime}|+1)+2$ (since every component in $H_{1}$ has a credit of 1, $\mathrm{cr}(B)=2$, and the newly formed component receives 2 credits). Plugging the upper bound on $\mathrm{cr}(H_{1})$ from above, we have $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H)-|K|+2-(|K^{\prime}|+1)+2=\mathrm{cr}(H)-|K|-|K^{\prime}|+3$. Thus, $\mathrm{cost}(H_{2})=|H_{2}|+\mathrm{cr}(H_{2})\leq|H|+|K|+|K^{\prime}|+\mathrm{cr}(H)-|K|-|K^{\prime}|+3=|H|+\mathrm{cr}(H)+3=\mathrm{cost}(H)+3$, which completes the proof. ∎
6.2 Proof of Lemma 10
@@
Let $C_{L}$ be the huge component such that the node $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
@@
Since $H\neq C_{L}$, ^$\bm{\hat{}}{G}_{H}$ has at least 2 nodes, and as ^$\bm{\hat{}}{G}_{H}$ is 2EC, $L$ has a neighbor, say $A$. By Proposition 8, we know that there must be a 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$. First, note that if $C_{A}$ is a large component, we simply add any two edges of $M_{3}$ to $H$ to obtain $H^{\prime}$, and it can be easily checked that $H^{\prime}$ meets our goal. Next assume that $C_{A}$ is a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. In either case, there must be two distinct edges in $M_{3}$, say $e_{1}$ and $e_{2}$, whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. We set $H^{\prime}\coloneq(H\setminus\{e\})\cup\{e_{1},e_{2}\}$ and observe that $H^{\prime}$ meets our goal. The only cases that remain are $C_{A}$ is a $\mathcal{C}_{6}$ or a $\mathcal{C}_{7}$. We consider the following exhaustive cases to complete the proof.
Case 1: There are two distinct edges in $M_{3}$, say $e_{1}$ and $e_{2}$, which are incident to $u$ and $v$ in $C_{A}$ and there is a Hamiltonian Path $P$ of $G[V(C_{A})]$ from $u$ to $v$. We set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup\{e_{1},e_{2}\}\cup P$ and, like before, it is easy to check that $H^{\prime}$ meets our goal.
@@
Case 2: $A$ is not a pendant node in ^$\bm{\hat{}}{G}_{H}$ (i.e., it has other neighbors other than $A$). Then, Lemma 3 is applicable, implying a 4-matching $M_{4}$ between $C_{A}$ and $C_{L}$. This is because we can partition $V(G)$ in two parts to have $V(C_{L})$ in one part and $V(C_{A})\cup V(C_{B})$ in another part (see Proposition 7) such that the only edges crossing the parts are between $V(C_{L})$ and $V(C_{A})$, and $|V(C_{A})|+|V(C_{B})|\geq 6+4=10$ and $|V(C_{L})|\geq 10$, because $L$ is huge. Arguing as before, there must exist edges $e_{1}$ and $e_{2}$ whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. Therefore, we conclude that actually Case 1 holds as $C_{A}\setminus\{e\}$ forms a Hamiltonian path of $G[V(C_{A})]$.
Case 3: $A$ is pendant node of ^$\bm{\hat{}}{G}_{H}$ (i.e., it has exactly one neighbor in ^$\bm{\hat{}}{G}_{H})$. Using the fact that we are not in Case 1 or Case 2, we prove that $C_{A}$ is contractible, which contradicts the fact that $G$ is structured, and hence Case 3 is redundant.
@@
To prove that $C_{A}$ is contractible, let the vertices of the cycle $C_{A}$ be labelled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{1}$ if $C_{A}$ is a $\mathcal{C}_{6}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{7}-v_{1}$ if $C_{A}$ is a $C_{7}$. Since we are not in Case 1 and there is the 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$, $C_{L}$ is adjacent to exactly three mutually non-adjacent vertices of $C_{A}$ via $M_{3}$, say w.l.o.g. $\{v_{1},v_{3},v_{5}\}$ (up to symmetry this is the only possibility in both the $\mathcal{C}_{6}$ and $\mathcal{C}_{7}$ cases). We say an edge in $G[V(C_{A})]$ is a nice diagonal if both its endpoints are in $\{v_{2},v_{4},v_{6}\}$. Therefore, the set of nice diagonals are $\{v_{2}v_{4},v_{4}v_{6},v_{6}v_{1}\}$. If there exists a nice diagonal $e$ in $G[V(C_{A})]$, then observe that there are two distinct edges in $M_{3}$ that are incident to say $u$ and $v$ in $C_{A}$ and there is a Hamiltonian Path $P\subseteq\{e\}\cup E(C_{A})$ from $u$ to $v$ in $G[V(C_{A})]$. For example, if the diagonal $v_{2}v_{4}$ exists in $G$ and $C_{A}$ is $\mathcal{C}_{6}$, then the desired Hamiltonian path is $v_{3}-v_{2}-v_{4}-v_{5}-v_{6}-v_{1}$ (for $\mathcal{C}_{6}$, up to symmetry, this is the only case; the $\mathcal{C}_{7}$ case is left to the reader to verify, a simple exercise). Hence, we are in fact in Case 1. Otherwise, if none of these nice diagonals are in $G$, observe that $C_{A}$ is contractible: The three sets of edges of $G$, namely, edges incident on $v_{2}$, edges incident on $v_{4}$, and edges incident on $v_{6}$ are all inside $G[V(C_{A})]$ and mutually disjoint. As any 2-ECSS of $G$ must contain at least $2$ edges incident on each vertex, it has to contain at least $6$ edges from $G[V(C_{A})]$, and hence $C_{A}$ is contractible. ∎
6.3 Proof of Lemma 11
@@
Let $C_{L}$ be a huge component and $C_{A}$ be a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$ component of $H$ such that $L$ and $A$ are in the same non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$. Given such an $L$, $A$, and $S$, we can find in polynomial time a cycle $K$ in ^$\bm{\hat{}}{G}_{H}$ through the nodes $L$ and $A$, where the two edges of the cycle $K$ entering the node $A$ are incident on two vertices $u$ and $v$ of $C_{A}$, such that at least one of the following two conditions holds.
@@
    We can find a node $D$ of ^$\bm{\hat{}}{G}_{H}$ outside of the cycle $K$ and a set of edges $F\subseteq E(G)$ disjoint to $E(K)$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
Using the above lemma, we prove Lemma 11, and then prove Lemma 12.
Proof of Lemma 11.
@@
We are given a non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$ which has a node $L$ such that $C_{L}$ is a huge component. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$. By definition, $S$ is a maximal 2-node-connected subgraph of ^$\bm{\hat{}}{G}_{H}$ such that each node is part of a cycle of length at least 3.
@@
If $S$ also contains a node $A$ corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$, then we apply Lemma 12 to obtain a cycle $K$ through $L$ and $A$ such that $K$ is incident to the vertices $u$ and $v$ of $C_{A}$. Now either Case (a) or Case (b) in the lemma statement holds.
@@
If Case (a) holds, then there is a Hamiltonian Path $P$ from $u$ to $v$ in $G[V(C_{A})]$ and we set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup K\cup P$. Notice that $H^{\prime}$ is a canonical bridgeless 2-edge cover and has fewer components compared to $H$; in $H^{\prime}$ we have created a cycle using $K$ and $P$ that passes through some 2EC components (including $L$) and all the vertices of $C_{A}$ using the Hamiltonian path $P$. Now, observe that $|H^{\prime}|\leq|H|+|K|-1$, and as each component of $H$ has a credit of at least 1, the large component corresponding to $L$ has credit $2$, and the newly created huge component in $H^{\prime}$ receives a credit of 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2)+2=\mathrm{cr}(H)-|K|+1$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|-1+\mathrm{cr}(H)-|K|+1=|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$.
@@
Otherwise, Case (b) holds, and there exist a node $D$ which is not in the cycle $K$, and a set of edges $F$ disjoint to $K$, such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})])$. Here, we set $H^{\prime}\coloneq(H\setminus(E(C_{A})\cup E(C_{D})))\cup K\cup F$. Like before, $H^{\prime}$ is bridgeless and has fewer components compared to $H$ for the following reason. From the condition of Case (b), the vertices of $C_{A}$ and $C_{D}$ remain 2EC in $H^{\prime}$ with the edges of $F$ and an additional edge $uv$; in $H^{\prime}$ the work of the edge $uv$ is instead done by a $u-v$ path induced by the cycle $K$. Also, the cycle $K$ 2-edge connects the various 2EC components (including $L$) with the vertices of $C_{A}$ and $C_{D}$ with the help of $F$. Thus, no bridges are introduced. Furthermore, observe that $|H^{\prime}|\leq|H|+|K|$ as $|F|=|E(C_{A})|+|E(C_{D})|$, and since each component in $H$ has at least a credit of 1, the huge component corresponding to $L$ has credit $2$, $C_{D}$ is not a component of $H^{\prime}$ and was not on $K$, and the newly created huge component receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2+1)+2=\mathrm{cr}(H)-|K|$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|=|H|+\mathrm{cr}(H)=\mathrm{cr}(H)$.
@@
Hence, we can assume that $S$ does not contain a node corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. Thus, all nodes of $S$ correspond to components that are either large, a $\mathcal{C}_{6}$ or a $\mathcal{C}_{7}$. Temporarily, we drop the parallel edges from the segment $S$ in the component graph ^$\bm{\hat{}}{G}_{H}$ keeping precisely one edge between adjacent nodes in $S$. Observe that $S$ continues to be 2-node connected even after dropping the parallel edges. Now, the smallest cycle through $L$ in $S$ in the component graph has length at least $3$ (as there are no parallel edges). Compute any cycle $K$ of length at least $3$ through $L$ in the segment $S$ and set $H^{\prime}\coloneq H\cup K$. $H^{\prime}$ is clearly a canonical bridgeless 2-edge cover of $G$ with fewer components. We have $|H^{\prime}|=|H|+|K|$ and since all components incident on the cycle $K$ have a credit of at least $\frac{3}{2}$ and $L$ has a credit of $2$, and they merge into a single huge component in $H^{\prime}$ which receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(\frac{3}{2}(|K|-1)+2)+2=\mathrm{cr}(H)-\frac{3}{2}|K|+\frac{3}{2}\leq\mathrm{cr}(H)-|K|$, as $|K|\geq 3$ implies $\frac{|K|}{2}\geq\frac{3}{2}$. Combining the above facts, we have $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|\leq|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$. ∎

### v2

@@ 6.1 Proof of Lemma˜8
Without loss of generality, $H$ has at least two components; if $H$ consists of a single component, it has to be huge as $\lvert V(H)\rvert=\lvert V(G)\rvert\geq\frac{4}{\varepsilon}\geq 10$. Let $A$ be a node in $\bm{\hat{}}{G}_{H}$. Let $K$ be some cycle in $\bm{\hat{}}{G}_{H}$ containing $A$, which must exist (and can be easily found) as $\bm{\hat{}}{G}_{H}$ is 2EC and has at least 2 nodes.
@@
Otherwise, arguing as before, we can find another cycle $K^{\prime}$ through $B$ in $\bm{\hat{}}{G}_{H_{1}}$ and add $K^{\prime}$ to $H_{1}$ to obtain $H_{2}$, i.e., $H_{2}\coloneq H\cup K\cup K^{\prime}$. Observe that $H_{2}$ is again a canonical bridgeless 2-edge cover and the newly formed component (containing the vertices of $B$) has at least 12 vertices, and is huge. We have $|H_{2}|=|H|+|K|+|K^{\prime}|$ and $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H_{1})-(|K^{\prime}|+1)+2$ (since every component in $H_{1}$ has a credit of at least 1, $\mathrm{cr}(B)=2$, and the newly formed component receives 2 credits). Plugging the upper bound on $\mathrm{cr}(H_{1})$ from above, we have $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H)-|K|+2-(|K^{\prime}|+1)+2=\mathrm{cr}(H)-|K|-|K^{\prime}|+3$. Thus, $\mathrm{cost}(H_{2})=|H_{2}|+\mathrm{cr}(H_{2})\leq|H|+|K|+|K^{\prime}|+\mathrm{cr}(H)-|K|-|K^{\prime}|+3=|H|+\mathrm{cr}(H)+3=\mathrm{cost}(H)+3$, which completes the proof. ∎
6.2 Proof of Lemma˜9
LLAABB
(a) $C_{A}$ is a $\mathcal{C}_{7}$ and there is a 4-matching between vertices of $C_{L}$ and $C_{A}$.
LLAA
(b) $C_{A}$ is a $\mathcal{C}_{7}$ and there is a nice diagonal in $C_{A}$.
LLAA
(c) $C_{A}$ is a $\mathcal{C}_{7}$ and $C_{A}$ is contractible as every solution must buy at least $6$ edges of $G[V(C_{A})]$ (e.g., the edges in blue).
Figure 1: Illustrations of different cases considered in the proof of Lemma˜9. Nodes of $\bm{\hat{}}{G}_{H}$ are depicted by a circle with black border, and vertices of $G$ are depicted by a black, filled circle. Edges of $H$ are black and solid, edges of $G\setminus H$ are gray and dashed, edges of $G\setminus H$ that are added to $H$ to obtain $H^{\prime}$ are red, and edges that are removed from $H$ to obtain $H^{\prime}$ are green.
@@ 6.2 Proof of Lemma˜9
Let $C_{L}$ be a huge component such that the node $L$ is a trivial segment of $\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
@@
Since $H\neq C_{L}$, $\bm{\hat{}}{G}_{H}$ has at least 2 nodes, and as $\bm{\hat{}}{G}_{H}$ is 2EC, $L$ has a neighbor, say $A$. By Proposition˜7, we know that there must be a 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$. First, note that if $C_{A}$ is a large component, we simply add any two edges of $M_{3}$ to $H$ to obtain $H^{\prime}$, and it can be easily checked that $H^{\prime}$ meets our goal as $\mathrm{cost}(H^{\prime})-\mathrm{cost}(H)=(|H^{\prime}|-|H|)+(\mathrm{cr}(H^{\prime})-\mathrm{cr}(H))\leq 2+(2-2-2)\leq 0$ (we add 2 edges, $\mathrm{cr}(C_{L})=\mathrm{cr}(C_{A})=2$, and, in $H^{\prime}$, $C_{L}$ and $C_{A}$ combine to form a single component of credit $2$). Thus, from now on, we assume $C_{A}$ is $\mathcal{C}_{i}$ for some $i\in\{4,5,6,7\}$.
Condition ($\ast$): If there are two distinct edges between $L$ and $A$, say $e_{1}$ and $e_{2}$, that are incident to two distinct vertices in $C_{A}$, say $u$ and $v$, such that there is a Hamiltonian Path $P$ of $G[V(C_{A})]$ from $u$ to $v$, then we set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup\{e_{1},e_{2}\}\cup P$, which is clearly a canonical bridgeless 2-edge-cover of $G$ with fewer components than $H$. Furthermore, $\mathrm{cost}(H^{\prime})-\mathrm{cost}(H)=(|H^{\prime}|-|H|)+(\mathrm{cr}(H^{\prime})-\mathrm{cr}(H))=(2-1)+(2-2-\mathrm{cr}(C_{A}))\leq 0$, since we added two edges $e_{1}$ and $e_{2}$, effectively removed 1 edge as $|C_{A}|-|P|=1$, $\mathrm{cr}(C_{L})=2$ and $\mathrm{cr}(C_{A})\geq 1$ in $H$, whereas in $H^{\prime}$ those combine to form a single component with credit 2. Thus, $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$, meeting our goal.
@@
We now show that Condition ($\ast$) always holds. If $C_{A}$ is a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$, then in either case, there must be two distinct edges in the matching $M_{3}$, say $e_{1}$ and $e_{2}$, whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. Thus, Condition ($\ast$) holds by setting $P=E(C_{A})\setminus\{e\}$. Otherwise, $C_{A}$ is a $\mathcal{C}_{6}$ or $\mathcal{C}_{7}$. If $A$ is not a pendant node in $\bm{\hat{}}{G}_{H}$ (i.e., it has other neighbors other than $A$), then, Lemma˜3 is applicable, implying a 4-matching $M_{4}$ between $C_{A}$ and $C_{L}$. This is because we can partition $V(G)$ in two parts to have $V(C_{L})$ in one part and $V(C_{A})\cup V(C_{B})$ in another part (see Proposition˜6) such that the only edges crossing the parts are between $V(C_{L})$ and $V(C_{A})$, and $|V(C_{A})|+|V(C_{B})|\geq 6+4=10$ and $|V(C_{L})|\geq 10$, because $L$ is huge. Arguing as before, there must exist edges $e_{1}$ and $e_{2}$ whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. Again, Condition ($\ast$) holds by setting $P=E(C_{A})\setminus\{e\}$. An illustration of this case is given in Figure˜1(a).
@@
Finally, we consider the case that $C_{A}$ is a $\mathcal{C}_{6}$ or $\mathcal{C}_{7}$ and $A$ is pendant node of $\bm{\hat{}}{G}_{H}$, i.e., the only neighbor it has is $L$. Assuming that Condition ($\ast$) does not hold, we prove that $C_{A}$ is contractible, which contradicts the fact that $G$ is structured. Let the vertices of the cycle $C_{A}$ be labeled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{1}$ if $C_{A}$ is a $\mathcal{C}_{6}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{7}-v_{1}$ if $C_{A}$ is a $C_{7}$. Since Condition ($\ast$) does not hold and there is the 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$, $C_{L}$ is adjacent to exactly three mutually non-adjacent vertices of $C_{A}$ via $M_{3}$, say w.l.o.g. $\{v_{1},v_{3},v_{5}\}$ (up to symmetry this is the only possibility in both the $\mathcal{C}_{6}$ and $\mathcal{C}_{7}$ cases). We say an edge in $G[V(C_{A})]$ is a nice diagonal if both its endpoints are in $\{v_{2},v_{4},v_{6}\}$. Therefore, the set of nice diagonals are $\{v_{2}v_{4},v_{4}v_{6},v_{6}v_{1}\}$. If there exists a nice diagonal $e$ in $G[V(C_{A})]$, then observe that there are two distinct edges in $M_{3}$ that are incident to say $u$ and $v$ in $C_{A}$, and there is a Hamiltonian Path $P\subseteq\{e\}\cup E(C_{A})$ from $u$ to $v$ in $G[V(C_{A})]$. For example, if the diagonal $v_{2}v_{4}$ exists in $G$ and $C_{A}$ is $\mathcal{C}_{6}$, then the desired Hamiltonian path is $v_{3}-v_{2}-v_{4}-v_{5}-v_{6}-v_{1}$ (for $\mathcal{C}_{6}$, up to symmetry, this is the only case; the $\mathcal{C}_{7}$ case is left to the reader to verify, a simple exercise). Hence, Condition ($\ast$) holds, a contradiction. An illustration of this case is given in Figure˜1(b). Otherwise, if none of these nice diagonals are in $G$, we can conclude that $C_{A}$ is contractible: The three sets of edges of $G$, namely, edges incident on $v_{2}$, edges incident on $v_{4}$, and edges incident on $v_{6}$ are all inside $G[V(C_{A})]$ and mutually disjoint. As any 2-ECSS of $G$ must contain at least $2$ edges incident on each vertex, it has to contain at least $6$ edges from $G[V(C_{A})]$, and hence $C_{A}$ is contractible. An illustration of this case is given in Figure˜1(c). ∎
6.3 Proof of Lemma˜10
@@
Let $S$ be a non-trivial segment of $\bm{\hat{}}{G}_{H}$ containing nodes $L$ and $A$ such that $C_{L}$ is huge and $C_{A}$ is a $\mathcal{C}_{4}$ or $\mathcal{C}_{5}$. We can compute in polynomial time a cycle $K$ in $\bm{\hat{}}{G}_{H}$ that passes through the nodes $L$ and $A$, where the two edges of the cycle $K$ entering the node $A$ are incident on two distinct vertices $u$ and $v$ of $C_{A}$ in $H$, and at least one of the following two conditions holds.
@@
    We can find a node $D$ of $\bm{\hat{}}{G}_{H}$ that is not in the cycle $K$ and a set of edges $F\subseteq E(G)$ disjoint to the edges of the cycle $K$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
Using the above lemma, we prove Lemma˜10, and then prove Lemma˜11.
Proof of Lemma˜10.
@@
We are given a non-trivial segment $S$ of $\bm{\hat{}}{G}_{H}$ which has a node $L$ such that $C_{L}$ is a huge component. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
@@
If $S$ also contains a node $A$ corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$, then we apply Lemma˜11 to obtain a cycle $K$ through $L$ and $A$ such that $K$ is incident to the vertices $u$ and $v$ of $C_{A}$. Now either Case (a) or Case (b) in the lemma statement holds.
@@
If Case (a) holds, then there is a Hamiltonian Path $P$ from $u$ to $v$ in $G[V(C_{A})]$, and we set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup K\cup P$. Notice that $H^{\prime}$ is a canonical bridgeless 2-edge cover and has fewer components compared to $H$; in $H^{\prime}$ we have created a cycle using $K$ and $P$ that passes through some 2EC components (including $L$) and all the vertices of $C_{A}$ using the Hamiltonian path $P$. Now, observe that $|H^{\prime}|\leq|H|+|K|-1$, and as each component of $H$ has a credit of at least 1, the large component corresponding to $L$ has credit $2$, and the newly created huge component in $H^{\prime}$ receives a credit of 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2)+2=\mathrm{cr}(H)-|K|+1$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|-1+\mathrm{cr}(H)-|K|+1=|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$.
@@
Otherwise, Case (b) holds, and there exist a node $D$ which is not in the cycle $K$, and a set of edges $F$ disjoint to $K$, such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$. Here, we set $H^{\prime}\coloneq(H\setminus(E(C_{A})\cup E(C_{D})))\cup K\cup F$. Like before, $H^{\prime}$ is bridgeless and has fewer components compared to $H$ for the following reason. From the condition of Case (b), the vertices of $C_{A}$ and $C_{D}$ remain 2EC in $H^{\prime}$ with the edges of $F$ and an additional edge $uv$; in $H^{\prime}$ the work of the edge $uv$ is instead done by a uvu-v path induced by the cycle $K$. Also, the cycle $K$ 2-edge connects the various 2EC components (including $L$) with the vertices of $C_{A}$ and $C_{D}$ with the help of $F$. Thus, no bridges are introduced. Furthermore, observe that $|H^{\prime}|\leq|H|+|K|$ as $|F|=|E(C_{A})|+|E(C_{D})|$, and since each component in $H$ has at least a credit of 1, the huge component corresponding to $L$ has credit $2$, $C_{D}$ is not a component of $H^{\prime}$ and was not on $K$, and the newly created huge component receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2+1)+2=\mathrm{cr}(H)-|K|$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|=|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$.
@@
Hence, we can assume that $S$ does not contain a node corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. Thus, all nodes of $S$ correspond to components that are either large, a $\mathcal{C}_{6}$ or a $\mathcal{C}_{7}$. Temporarily, we drop the parallel edges from the segment $S$ in the component graph $\bm{\hat{}}{G}_{H}$ keeping precisely one edge between adjacent nodes in $S$. Observe that $S$ continues to be 2-node connected even after dropping the parallel edges. Now, the smallest cycle through $L$ in $S$ in the component graph has length at least $3$ as there are no parallel edges. Compute any cycle $K$ of length at least $3$ through $L$ in the segment $S$ and set $^{\prime}\coloneq H\cup K. HH^{\prime}$ is clearly a canonical bridgeless 2-edge cover of $G$ with fewer components. We have $|H^{\prime}|=|H|+|K|$ and since all components incident on the cycle $K$ have a credit of at least $\frac{3}{2}$ and $L$ has a credit of $2$, and they merge into a single huge component in $H^{\prime}$ which receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(\frac{3}{2}(|K|-1)+2)+2=\mathrm{cr}(H)-\frac{3}{2}|K|+\frac{3}{2}\leq\mathrm{cr}(H)-|K|$, as $|K|\geq 3$ implies $\frac{|K|}{2}\geq\frac{3}{2}$. Combining the above facts, we have $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|\leq|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$. ∎